find the integral solution of the equation given below.  please give me the answer also with propar
solution.???
           
           x^2-13y^2=(-1).

Grade:11

1 Answers

SAHIL
7 years ago
= 02 + -13y2 = 01 + X2 + -13y2Combine like terms: 13y2 + -13y2 = 13y2 + -13y21 + X2 + -13y2 = 0 + 13y2 + -13y2Combine like terms: -1 + 1 = 01 + X2 + -13y2 = -1 + 1 + 13y2 + -13y2Reorder the terms:1 + X2 + 1 + -13y2 = -1 + 13y2 + -13y2Reorder the terms:1 + X2 = -1 + 13y2SimplifyingX2 = -1 + 13y2X2 + 0 = -1 + 13y2 = 0X2 + 13y2Combine like terms: -13y2 = -1 + 13y2 + 13y2 + -13y2' to each side of the equation.X2 = -1Solving for variable 'X'.Move all terms containing X to the left, all other terms to the right.Add '13y2 + -13y2 = -1SolvingX2 + -13y2SimplifyingXHENCE SOLUTION TO THIS CANNOT BE DETERMINED